OpenStudy (shelby1290):

Solve by substitution: x+3y=0 3x-6y=5

OpenStudy (anonymous):

multiply the first equation straight across by 2, then add, the y terms will go bye bye

OpenStudy (misty1212):

HI!!

OpenStudy (shelby1290):

ok so that would be 2x+6y=0

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

then add in a column

OpenStudy (shelby1290):

Hi

OpenStudy (shelby1290):

Why do i need to add in a column?

OpenStudy (misty1212):

if you do it you will see why

OpenStudy (misty1212):

what @satellite73 said, the y terms will drop out because \(6y-6y=0\)

OpenStudy (shelby1290):

How do you add in a column....?

OpenStudy (misty1212):

\[2x+6y=0\\ 3x-6y=5\] andd and get \[5x=5\]

OpenStudy (triciaal):

@satellite73 It said by substitution

OpenStudy (anonymous):

there is more than one way to skin a cat

OpenStudy (xapproachesinfinity):

just do it this way lol

OpenStudy (misty1212):

it should be clear what \(x\) is if \(5x=5\) right?

OpenStudy (shelby1290):

lol that saying though @satellite73

OpenStudy (triciaal):

there are 4 methods used to solve simultaneous equations

OpenStudy (shelby1290):

x=1

OpenStudy (misty1212):

yes

OpenStudy (triciaal):

the directions specifically said to solve by substitution

OpenStudy (misty1212):

then if \(x=1\) and \(x+3y=0\) that means \[1+3y=0\] and you can solve that one for \(y\)

OpenStudy (shelby1290):

@triciaal what are we doing incorrectly?

OpenStudy (triciaal):

"Solve by substitution:" @ shelby1290 not following the directions

OpenStudy (triciaal):

@shelby1290 the question said what method to use

OpenStudy (shelby1290):

bring 1 to the other side.... 3y=0-1 3y=1 y=0.333333 lol

OpenStudy (misty1212):

i would probably say \(-\frac{1}{3}\) dont forget the minus sing

OpenStudy (shelby1290):

@triciaal I've checked my textbook and the answer is x=1 and y= -1/3

OpenStudy (shelby1290):

yeaaa

OpenStudy (misty1212):

lol "sign"

OpenStudy (triciaal):

whatever method you use will give the same answer

OpenStudy (misty1212):

or the minus sing too, an nice tune

OpenStudy (shelby1290):

lol sign*

OpenStudy (shelby1290):

But thank you mistyyyy

OpenStudy (misty1212):

*sigh...

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (triciaal):

@shelby1290 as I said before there are 4 different methods that can be used to solve these equations

OpenStudy (shelby1290):

Yes I understand that.

OpenStudy (triciaal):

the is the addition / elimination method where you eliminate one variable when you add the equations

OpenStudy (shelby1290):

Thank you for letting me know

OpenStudy (triciaal):

to use the substitution method let one variable be the subject and use this expression for the variable in the second equation.

OpenStudy (triciaal):

for example let x = -3y from the 1st equation then use this for x in the 2nd 3x - 6y = 5 3(-3y) -6y = 5 -15y = 5 so y = -1/3 use this value in the expression to get x x = -3(-1/3) = -1 different method same answer (-1, -1/3)